# Position: LLMs Can be Good Tutors in Foreign Language Education

### v1

@@ Abstract
While recent efforts have begun integrating large language models (LLMs) into foreign language education (FLE), they often rely on traditional approaches to learning tasks without fully embracing educational methodologies, thus lacking adaptability to language learning. To address this gap, we argue that LLMs have the potential to serve as effective tutors in FLE. Specifically, LLMs can play three critical roles: (1) as data enhancers, improving the creation of learning materials or serving as student simulations; (2) as task predictors, serving as learner assessment or optimizing learning pathway; and (3) as agents, enabling personalized and inclusive education. We encourage interdisciplinary research to explore these roles, fostering innovation while addressing challenges and risks, ultimately advancing FLE through the thoughtful integration of LLMs.
Machine Learning, ICML
@@ 1 Introduction
Foreign Language Education (FLE) has long been a cornerstone of global education and a critical component of K-12 curricula, equipping students with the linguistic and cultural competencies necessary for an interconnected world. However, traditional FLE methods often fall short in addressing the diverse needs of learners (Hou, 2020). Challenges such as limited personalization, scalability constraints, and the lack of real-time feedback are particularly pronounced in large classroom settings (Ehrenberg et al., 2001), where individual attention is scarce, leaving many students disengaged or struggling to keep pace. Addressing these shortcomings requires innovative approaches that not only enhance the quality of instruction but also adapt to the unique learning trajectories of students (Eaton, 2010).
@@
In parallel, the advent of LLMs has opened new possibilities for transforming FLE (Caines et al., 2023). These models, powered by large-scale pre-training, exhibit remarkable capabilities in natural language understanding and generation, making them promising candidates for roles traditionally filled by human tutors. By leveraging LLMs as AI tutors, we may overcome many of the inherent limitations of conventional teaching methods, offering scalable, interactive, and personalized learning experiences (Chen et al., 2024; Schmucker et al., 2024). Therefore, this position paper argues that LLMs can be effective tutors in foreign language education, complementing human expertise and addressing key limitations of traditional methods.
Figure 1: Involved disciplines of LLM for FLE.
@@
Figure 2: Overview of roles of LLMs in foreign language education. An overview of related literature is provided in Appendix A.
@@
As shown in Figure 1, FLE intersects with multiple disciplines, each of which underscores the potential of LLMs to revolutionize this domain. From the perspective of (1) computer science, advancements in machine learning and NLP have enabled LLMs to process and generate human-like language at an unprecedented scale; (2) linguistics (Radford et al., 2009) contributes a deeper understanding of grammar, phonetics, and semantics, allowing LLMs to generate accurate and understandable language outputs; (3) education provides the foundation for designing effective pedagogical strategies, ensuring that LLMs can deliver personalized, engaging, and developmentally appropriate learning experiences; and finally, (4) psycholinguistics (Steinberg & Sciarini, 2013) bridges the gap between language acquisition and cognitive processes, enabling LLMs to optimize learner interactions by adapting to individual needs and fostering meaningful engagement. Together, these disciplines position LLMs as uniquely capable of addressing the multifaceted challenges of foreign language education.
@@
Moreover, FLE encompasses four core skills: listening, speaking, reading, and writing, each of which can be significantly enhanced by LLMs. For listening, LLMs can generate diverse audio materials (Ghosal et al., 2023) and facilitate interactive voice-based exercises, helping learners improve their ability to discern pronunciation, intonation, and contextual meaning. In speaking, LLMs can simulate realistic conversations (Siyan et al., 2024), provide pronunciation feedback, and scaffold learners’ oral communication skills through iterative practice. For reading, LLMs can curate leveled texts, generate comprehension questions (Samuel et al., 2024), and engage learners in discussions that deepen their understanding of written content. Finally, in writing, LLMs can offer real-time grammar, syntax, and style feedback while assisting with idea generation and iterative revisions (Stahl et al., 2024a). By addressing these core skills holistically, LLMs have the potential to deliver a comprehensive and adaptive learning experience.
@@
Despite these opportunities, the deployment of LLMs in FLE must be approached carefully, ensuring that their integration complements rather than replaces human tutors (Jeon & Lee, 2023). As illustrated in Figure 2, this paper explores three critical roles of LLMs in this context: their function as data enhancers (Section 4) to optimize learning materials, their capacity as task predictors (Section 5) to tailor educational solutions, and their potential as agents (Section 6) that deliver interactive and adaptive language instruction. By examining these roles, we aim to demonstrate how LLMs can address the limitations of traditional FLE methods while advancing our understanding of intelligent tutoring systems. Additionally, we discuss potential challenges, ethical considerations, and future directions (Section 7) for integrating LLMs into FLE, offering a technical guideline for researchers and educators to harness their transformative potential. We also describe the paradigm shift of leveraging AI for FLE, starting from the last century, as one of our contributions in Section 3.
Figure 3: Roadmap of Foreign Language Education.
@@ 2.1 Foreign Language Education
FLE (Watzke, 2003; Rashov, 2024) has long been a cornerstone of global communication and cultural exchange. Traditional methods often emphasize grammar rules, vocabulary memorization, and repetitive practice, supplemented by limited opportunities for real-world application. Such approaches are often constrained by the availability of skilled teachers, the diversity of learners’ needs, and the lack of personalized feedback (Williams et al., 2004). To bridge the gaps, many technologies for FLE have been proposed (Alhusaiyan, 2024), focusing on solving specific tasks instead of describing the whole picture of foreign language tutoring. While intelligent language tutoring systems (Schwind, 1990; C Angelides & Garcia, 1993) have the potential to create adaptive environments, attention to this field is relatively less compared to other subjects like science and mathematics. One key reason lies in the inherent complexity of language as an ill-defined domain (Schmidt & Strasser, 2022), posing a great challenge in establishing a valid automatic analysis of learner languages due to the vast variability and unpredictability of human language.
2.2 Large Language Models
The potential of LLMs in education (Alhafni et al., 2024), particularly in FLE (Gao et al., 2024; Karataş et al., 2024), is immense. Benefiting from large-scale pre-training on extensive corpora, LLMs have demonstrated emergent abilities including (1) in-context learning (Dong et al., 2022), which allows the model to adapt to new tasks and provide contextually relevant responses based on a few examples provided during the interaction; (2) instruction following (Zeng et al., 2024), which enables the model to process and execute complex user instructions with high accuracy; and (3) reasoning and planning (Huang et al., 2024b), which allows the model to generate coherent, structured, and context-aware outputs, even for tasks that require multi-step thinking.
@@ 2.2 Large Language Models
However, these fundamental capabilities, while impressive, are not sufficient to fully meet the unique demands of FLE. Teaching a foreign language requires more than generating grammatically correct sentences or providing accurate translations; it demands a nuanced understanding of pedagogy, learner psychology, and cultural context. Maurya et al. (2024) propose an evaluation taxonomy that identifies eight critical dimensions for assessing AI tutors. These dimensions can be broadly categorized into two groups. (1) Problem-solving abilities assess the technical capabilities of LLMs to perform tasks relevant to FLE. (2) Pedagogical alignment abilities evaluate how well the LLM aligns with effective teaching and learning principles. Pedagogical alignment includes the model’s ability to adapt to the learner’s proficiency level, provide scaffolded feedback, foster engagement, and maintain motivation. While LLMs can give direct answers, their ability to replicate these nuanced teaching strategies remains a challenge (Wang et al., 2024a). Additionally, pedagogical alignment requires sensitivity to cultural and contextual factors, ensuring that the learning experience is relevant and meaningful to the student.
@@ 3 Paradigm Shift
The development of AI models for FLE can be broadly categorized into four successive generations as shown in Figure 3: (1) rule-based models, (2) statistical models, (3) neural models, and (4) large language models. We leave the detailed description in Appendix B.
Our position. We foresee next-generation LLMs with deeper alignment to pedagogical principles and stronger guardrails to mitigate misinformation and bias. Future models may integrate multimodal data (e.g., text, image, video, speech) and employ meta-learning to adapt in real time to diverse learner profiles. These improvements will reinforce the position that LLMs can evolve into even more effective tutors for FLE.
@@ 4 LLMs as Data Enhancers
Education is a high-stake area where any form of LLMs’ hallucination could cause devastating harm to humans’ cognition activities (Ho et al., 2024). One of the hallucination causes is from data (Huang et al., 2023). Therefore, high-quality and diverse data resources (Long et al., 2024) are critical to ensuring the reliability of incorporating LLMs into FLE. The 1) creation, 2) reformation, and 3) annotation of educational materials are crucial to delivering effective and engaging teaching. Traditional resource development methods often lack the scalability, adaptability, and personalization necessary to meet the diverse needs of learners (Feng et al., 2021; Shorten et al., 2021). In contrast, LLMs emerge as transformative tools capable of enhancing these processes (Wang et al., 2024c; Liu et al., 2024c). This section explores how LLMs serve as data enhancers in FLE.
@@ 4.1 Data Creation
Creating pedagogically sound and learner-specific data is a cornerstone of personalized learning. However, manually creating such resources is time-consuming and often fails to address the wide range of learner needs (Cochran et al., 2022). LLMs can revolutionize this process by generating tailored and diverse educational content or responses on demand (Zha et al., 2023; Cochran et al., 2023).
@@
A primary use of LLMs in data creation is the generation of educational questions aligned with specific learning objectives. Due to their superior contextual understanding, classic rule-based approaches have largely been eclipsed by neural network-based techniques (Kurdi et al., 2020; Rathod et al., 2022; Mulla & Gharpure, 2023). LLMs can produce answer-aware (whose target answer is known) or answer-agnostic (whose answer is open) (Zhang et al., 2021), resulting in more nuanced exercises and assessments (Xiao et al., 2023).
@@
Simulating the learner’s perspective is crucial for designing adaptive instructional materials. Traditional surveys and standardized tests often fail to capture the complexity of dynamic learner behaviors (Käser & Alexandron, 2024). In contrast, LLM-based approaches enable high-fidelity, context-aware student simulations (Liu et al., 2024d; Yue et al., 2024), generating synthetic learners who exhibit realistic mastery levels and evolving behaviors. For instance, Generative Students (Lu & Wang, 2024) create simulated learners with various competency levels, while EduAgent (Xu et al., 2024) integrates cognitive priors to model complex learning trajectories and behaviors better.
@@ Discussion.
While LLMs excel at generating educational content, current approaches mainly focus on question creation, leaving many areas of FLE underexplored. Essential tasks like generating culturally rich reading materials, context-dependent writing prompts, or dynamic comprehension exercises are still lacking in diversity and depth. Additionally, the student simulations created by LLMs often fail to reflect long-term learning trajectories or the intricacies of individual learning progress, leading to challenges in creating truly adaptive systems.
@@ 4.2 Data Reformation
Transforming existing materials into different forms can yield more comprehensive and immersive learning experiences. For example, Book2Dial(Wang et al., 2024b) generates teacher-student dialogues grounded in textbooks, keeping the content both relevant and informative. Their approach includes multi-turn question generation and answering(Kim et al., 2022), dialogue inpainting (Dai et al., 2022), and role-playing. Likewise, Slide2Lecture (Zhang-Li et al., 2024) automatically converts lecture slides into structured teaching agendas, enabling interactive follow-up and deeper learner engagement.
@@
Another vital application is simplifying or paraphrasing complex texts to specified readability levels (Huang et al., 2024a) without losing key concepts (Al-Thanyyan & Azmi, 2021). This is particularly beneficial in FLE settings, where language beginners often face advanced vocabulary and complex structures (Day et al., 2025). Recent advancements in controllable generation (Zhang et al., 2023) leverage model fine-tuning on curated datasets (Zeng et al., 2023) or decoding-time interventions (Liang et al., 2024), thereby allowing educators to specify text complexity, style, or tone.
@@
Beyond linguistic correctness, cultural nuance is pivotal in FLE (Byram, 1989, 2008). LLMs can facilitate this process by recontextualizing existing materials to reflect the cultural and social norms of the target language (Liu et al., 2024a). For instance, a short story originally set in an English-speaking environment may be adapted for French or Japanese learners by adjusting the characters’ names, idiomatic expressions, or social customs, while still preserving core instructional goals. This cultural adaptation not only enhances learner engagement but also strengthens cross-cultural competencies.

### v2

# Position: LLMs Can be Good Tutors in English Education

@@ Abstract
While recent efforts have begun integrating large language models (LLMs) into English education, they often rely on traditional approaches to learning tasks without fully embracing educational methodologies, thus lacking adaptability to language learning. To address this gap, we argue that LLMs have the potential to serve as effective tutors in English Education. Specifically, LLMs can play three critical roles: (1) as data enhancers, improving the creation of learning materials or serving as student simulations; (2) as task predictors, serving as learner assessment or optimizing learning pathway; and (3) as agents, enabling personalized and inclusive education. We encourage interdisciplinary research to explore these roles, fostering innovation while addressing challenges and risks, ultimately advancing English Education through the thoughtful integration of LLMs.
Position: LLMs Can be Good Tutors in English Education
Jingheng Ye1,2, Shen Wang1, Deqing Zou2, Yibo Yan1,3, Kun Wang1, Hai-Tao Zheng2,4thanks: Corresponding Author., Ruitong Liu2, Zenglin Xu5, Irwin King6, Philip S. Yu7, Qingsong Wen [^1] footnotemark: 1 1Squirrel Ai Learning, 2Tsinghua University, 3HKUST (Guangzhou), 4Peng Cheng Laboratory, 5Fudan University, 6The Chinese University of Hong Kong, 7University of Illinois at Chicago yejh22@mails.tsinghua.edu.cn
@@ 1 Introduction
English Education has long been a cornerstone of global education and a critical component of K-12 curricula, equipping students with the linguistic and cultural competencies necessary for an interconnected world Alhusaiyan (2025); Katinskaia (2025). However, traditional English teaching methods often fall short in addressing the diverse needs of learners Hou (2020). Challenges such as limited personalization, scalability constraints, and the lack of real-time feedback are particularly pronounced in large classroom settings Ehrenberg et al. (2001). Addressing these shortcomings requires innovative approaches that not only enhance the quality of instruction but also adapt to the unique learning trajectories of students Eaton (2010).
@@
Recently, LLMs have opened new possibilities for transforming English Education Caines et al. (2023). LLMs exhibit remarkable natural language understanding and generation capabilities, making them promising candidates for roles traditionally filled by human tutors. Leveraging LLMs as AI tutors can overcome the inherent limitations of conventional teaching methods, offering scalable, interactive, and personalized learning experiences Chen et al. (2024); Schmucker et al. (2024). Therefore, this position paper argues that LLMs can be effective tutors in English education, complementing human expertise and addressing key limitations of traditional methods.
Figure 1: Involved disciplines of LLM for English Edu.
@@
Figure 2: Overview of three roles of LLMs in English education. An overview of related literature is provided in Appendix A.
@@
As shown in Figure 1, English Education intersects with multiple disciplines, each of which underscores the potential of LLMs to revolutionize this domain. From the perspective of (1) computer science, advancements in machine learning and NLP have enabled LLMs to process and generate human-like language at an unprecedented scale; (2) linguistics Radford et al. (2009) contributes a deeper understanding of grammar, phonetics, and semantics, allowing LLMs to generate accurate and understandable language outputs; (3) education provides the foundation for designing effective pedagogical strategies, ensuring that LLMs can deliver personalized, engaging, and developmentally appropriate learning experiences; and finally, (4) psycholinguistics Steinberg and Sciarini (2013) bridges the gap between language acquisition and cognitive processes, enabling LLMs to optimize learner interactions by adapting to individual needs and fostering meaningful engagement. Together, these disciplines position LLMs as uniquely capable of addressing the multifaceted challenges of English education.
@@
Moreover, English Education encompasses four core skills: listening, speaking, reading, and writing, each of which can be significantly enhanced by LLMs. For listening, LLMs can generate diverse audio materials Ghosal et al. (2023) and facilitate interactive voice-based exercises, helping learners improve their ability to discern pronunciation, intonation, and contextual meaning. In speaking, LLMs can simulate realistic conversations Siyan et al. (2024), provide pronunciation feedback, and scaffold learners’ oral communication skills through iterative practice. For reading, LLMs can curate leveled texts, generate comprehension questions Samuel et al. (2024), and engage learners in discussions that deepen their understanding of written content. Finally, in writing, LLMs can offer real-time grammar, syntax, and style feedback while assisting with idea generation and iterative revisions Stahl et al. (2024a). By addressing these core skills holistically, LLMs have the potential to deliver a comprehensive and adaptive learning experience.
@@
Despite these opportunities, the deployment of LLMs in English Education must be approached carefully, ensuring that their integration complements rather than replaces human tutors Jeon and Lee (2023). As illustrated in Figure 2, this paper explores three critical roles of LLMs in this context: their function as data enhancers (Section 4) to optimize learning materials, their capacity as task predictors (Section 5) to tailor educational solutions, and their potential as agents (Section 6) that deliver interactive and adaptive language instruction. By examining these roles, we aim to demonstrate how LLMs can address the limitations of traditional English teaching methods while advancing our understanding of intelligent tutoring systems. Additionally, we discuss potential challenges (Section 7) and future directions (Appendix B) for integrating LLMs into English Education, offering a technical guideline for researchers and educators to harness their transformative potential. We also describe the paradigm shift of leveraging AI for English Education, starting from the last century, as one of our contributions in Section 3.
@@ 2.1 English Education
Traditional English Education methods often emphasize grammar rules, vocabulary memorization, and repetitive practice, supplemented by limited opportunities for real-world application Watzke (2003). Such approaches are often constrained by the availability of skilled teachers, the diversity of learners’ needs, and the lack of personalized feedback Williams et al. (2004). Recently, many technologies for English Education have been proposed Alhusaiyan (2024), focusing on solving specific tasks instead of describing the whole picture of English tutoring. While intelligent language tutoring systems have the potential to create adaptive environments, attention to this field is relatively less compared to other subjects like science Shao et al. (2025) and mathematics Ahn et al. (2024). One key reason lies in the inherent complexity of language as an ill-defined domain Schmidt and Strasser (2022), posing a great challenge in establishing a valid automatic analysis of learner languages due to the vast variability and unpredictability of human language.
2.2 Large Language Models for Education
@@
The potential of LLMs in education Alhafni et al. (2024), particularly in English Education Gao et al. (2024); Karataş et al. (2024); Cherednichenko et al. (2024), is immense. Benefiting from large-scale pre-training on extensive corpora, LLMs have demonstrated emergent abilities including (1) in-context learning Dong et al. (2022), which allows the model to adapt to new tasks and provide contextually relevant responses based on a few examples provided during the interaction; (2) instruction following Zeng et al. (2024), which enables the model to process and execute complex user instructions with high accuracy; and (3) reasoning and planning Huang et al. (2024b), which allows the model to generate coherent, structured, and context-aware outputs, even for tasks that require multi-step thinking. However, these fundamental capabilities, while impressive, are insufficient to fully meet the unique demands of English Education. Teaching English requires more than generating grammatically correct sentences or providing accurate translations; it demands a nuanced understanding of pedagogy, learner psychology, and cultural context. Maurya et al. (2024) propose an evaluation taxonomy that identifies eight critical dimensions for assessing AI tutors. These dimensions can be broadly categorized into two groups. (1) Problem-solving abilities assess the technical capabilities of LLMs to perform tasks relevant to English Education. (2) Pedagogical alignment abilities evaluate how well the LLM aligns with effective teaching and learning principles. Pedagogical alignment includes the model’s ability to adapt to the learner’s proficiency level, provide scaffolded feedback, foster engagement, and maintain motivation. While LLMs can give direct answers, their ability to replicate these nuanced teaching strategies remains a challenge Wang et al. (2024a).
@@ 3 Paradigm Shift
The development of AI models for English Education can be broadly categorized into four successive generations as shown in Figure 3: (1) rule-based models, (2) statistical models, (3) neural models, and (4) large language models.
Stage 1: Rule-based Models (1960s–1990s).
Early solutions relied on handcrafted linguistic rules to process language in tightly constrained scenarios Grosan et al. (2011); C Angelides and Garcia (1993). Classical platforms like PLATO Hart (1981) and Systran Toma (1977) operated effectively for highly structured tasks (e.g., grammar drills) but struggled with complex, context-dependent interactions.
Stage 2: Statistical Models (1990s–2010s).
With the increased availability of digitized corpora, methods such as the early version of Google Translate Och (2006) and Dragon NaturallySpeaking Blair (1997) pioneered statistical pattern mining. These approaches leveraged large datasets to infer linguistic rules and conduct specific tasks probabilistically, improving scalability yet still lacking deeper semantic understanding.
Stage 3: Neural Models (2010s–2020s).
The advent of deep learning architectures (e.g., RNNs Yu et al. (2019) and Transformers Vaswani (2017)) enabled more robust context modeling, sparking transformative applications like Grammarly Fitria (2021) and Duolingo Vesselinov and Grego (2012). These systems offered enhanced personalization and feedback, significantly augmenting learners’ writing and reading comprehension.
Stage 4: Large Language Models (2020s–Present).
Nowadays, various LLMs (e.g., ChatGPT Achiam et al. (2023)) combine massive pre-training with instruction tuning, achieving impressive results in multi-turn dialogue, individualized scaffolding, and multimodal integration. Tools such as Khanmigo Anand (2023) demonstrate LLMs’ potential for real-time conversational practice, dynamic content creation, and inclusive educational support at scale.
Our position. We foresee next-generation LLMs with deeper alignment to pedagogical principles and stronger guardrails to mitigate misinformation and bias. Future models may integrate multimodal data (e.g., text, image, video, speech) to adapt to diverse learner profiles in real time. These improvements will reinforce the position that LLMs can evolve into more effective tutors for English Education.
Figure 3: Roadmap of English Education.
@@ 4 LLMs as Data Enhancers
Education is a high-stake area where any hallucination could cause devastating harm to humans’ cognition activities Ho et al. (2024). One of the hallucination causes is from data Huang et al. (2023). Therefore, high-quality and diverse data resources Long et al. (2024) are critical to ensuring the reliability of incorporating LLMs into English Education. The 1) creation, 2) reformation, and 3) annotation of educational materials are crucial to delivering effective and engaging teaching. Traditional resource development methods often lack the scalability, adaptability, and personalization necessary to meet the diverse needs of learners Feng et al. (2021); Shorten et al. (2021). In contrast, LLMs emerge as transformative tools capable of enhancing these processes Wang et al. (2024c); Liu et al. (2024c). This section explores how LLMs serve as data enhancers in English Education.
@@ 4.1 Data Creation
Creating pedagogically sound and learner-specific data is a cornerstone of personalized learning. However, manually creating such resources is time-consuming and often fails to address the wide range of learner needs Cochran et al. (2022). LLMs can revolutionize this process by generating tailored and diverse educational content or responses on demand Zha et al. (2023); Cochran et al. (2023).
@@
A primary use of LLMs in data creation is the generation of educational questions aligned with specific learning objectives. Due to their superior contextual understanding, classic rule-based approaches have largely been eclipsed by neural network-based techniques Kurdi et al. (2020); Rathod et al. (2022); Mulla and Gharpure (2023). LLMs can produce answer-aware (whose target answer is known) or answer-agnostic (whose answer is open) Zhang et al. (2021), resulting in more nuanced exercises and assessments Xiao et al. (2023).
@@
Simulating the learner’s perspective is crucial for designing adaptive instructional materials. Traditional surveys and standardized tests often fail to capture the complexity of dynamic learner behaviors Käser and Alexandron (2024). In contrast, LLM-based approaches enable high-fidelity, context-aware student simulations Liu et al. (2024d); Yue et al. (2024), generating synthetic learners who exhibit realistic mastery levels and evolving behaviors. For instance, Generative Students Lu and Wang (2024) create simulated learners with various competency levels, while EduAgent Xu et al. (2024) integrates cognitive priors to model complex learning trajectories and behaviors better.
@@ Discussion.
While LLMs excel at generating educational content, current approaches mainly focus on question creation, leaving many areas of English Education underexplored. Essential tasks like generating culturally rich reading materials, context-dependent writing prompts, or dynamic comprehension exercises still lack diversity and depth. Additionally, the student simulations created by LLMs often fail to reflect long-term learning trajectories or the intricacies of individual learning progress.
@@ 4.2 Data Reformation
Transforming existing materials into different forms can yield more comprehensive and immersive learning experiences. For example, Book2DialWang et al. (2024b) generates teacher-student dialogues grounded in textbooks, keeping the content both relevant and informative. Their approach includes multi-turn question generation and answeringKim et al. (2022), dialogue inpainting Dai et al. (2022), and role-playing. Likewise, Slide2Lecture Zhang-Li et al. (2024) automatically converts lecture slides into structured teaching agendas, enabling interactive follow-up and deeper learner engagement.
@@
Another vital application is simplifying or paraphrasing complex texts to specified readability levels Huang et al. (2024a) without losing key concepts Al-Thanyyan and Azmi (2021). This is particularly beneficial in English Education settings, where language beginners often face advanced vocabulary and complex structures Day et al. (2025). Recent advancements in controllable generation Zhang et al. (2023) leverage model fine-tuning on curated datasets Zeng et al. (2023) or decoding-time interventions Liang et al. (2024), thereby allowing educators to specify text complexity, style, or tone.
@@
Beyond linguistic correctness, cultural nuance is another crucial factor in English Education Byram (1989, 2008). LLMs can facilitate this process by recontextualizing existing materials to reflect the cultural and social norms of different areas Liu et al. (2024a); Adilazuarda et al. (2024); Kharchenko et al. (2024). For instance, a short story originally set in an English-speaking environment may be adapted for Japanese students by adjusting the characters’ names, idiomatic expressions, or social customs, while preserving core instructional goals. This cultural adaptation not only enhances learner engagement but also strengthens cross-cultural competencies.